# Visual-Friendly Concept Protection via Selective Adversarial Perturbations

### v2

@@ Abstract
Personalized concept generation by tuning diffusion models with a few images raises potential legal and ethical concerns regarding privacy and intellectual property rights. Researchers attempt to prevent malicious personalization using adversarial perturbations. However, previous efforts have mainly focused on the effectiveness of protection while neglecting the visibility of perturbations. They utilize global adversarial perturbations, which introduce noticeable alterations to original images and significantly degrade visual quality. In this work, we propose the Visual-Friendly Concept Protection (VCPro) framework, which prioritizes the protection of key concepts chosen by the image owner through adversarial perturbations with lower perceptibility. To ensure these perturbations are as inconspicuous as possible, we introduce a relaxed optimization objective to identify the least perceptible yet effective adversarial perturbations, solved using the Lagrangian multiplier method. Qualitative and quantitative experiments validate that VCPro achieves a better trade-off between the visibility of perturbations and protection effectiveness, effectively prioritizing the protection of target concepts in images with less perceptible perturbations. Our code is available at https://github.com/KululuMi/VCPro.
Index Terms:
Unauthorized personalizations, Adversarial Attacks, Diffusions, Privacy protections, Copyright protections.
I Introduction
@@
With the advent of popular image generative models [1, 2, 3] such as Stable Diffusion [2] and GPT-4o [4], people lacking expertise in drawing or photography can effortlessly create realistic or artistic works using simple textual descriptions. However, the success of these models has raised significant concerns about privacy, intellectual property rights, and various legal and ethical issues [5]. For example, an adversary could easily generate sensitive specific concepts, such as personal fake images, or imitate renowned artworks for commercial purposes, using only a few reference images and concept personalization techniques such as Textual Inversion [6] or DreamBooth [7].
@@
Figure 1: The first row shows protected images from Anti-DreamBooth [8], Mist [9], MetaCloak [10], SDS(-) [11], PhotoGuard [12], and our VCPro. The second row highlights magnified background and face regions, while the third row presents Textual Inversion results. With 8/255 perturbation size, our method achieves a better balance between protection effectiveness and visual quality for user-important concepts, as shown by higher PSNR/SSIM values.
@@
One prevailing direction to mitigate these potential risks is to leverage adversarial attack techniques, transforming original images into adversarial examples, termed “protected images”. These protected images can misguide the personalized diffusion model and deceive its generation process, which resists malicious editing or personalization [12, 13, 14, 8, 15, 10]. For instance, AdvDM [13] employs adversarial attacks in the inversion stage of Stable Diffusion to safeguard against malicious imitation of the style of a specific artist. Later, its updated version, Mist [14], enhances the protection efficacy of protected images by adding a textual loss, extending its application from Textual Inversion to DreamBooth. Unlike AdvDM and Mist, which focus on art style protection, Anti-DreamBooth [8] undermines DreamBooth model generation quality to enhance privacy by adding adversarial perturbations to human face images before posting online. Furthermore, MetaCloak [10] leverages a meta-learning strategy and data transformations to generate more effective and robust protected images against DreamBooth.
@@
However, these methods primarily focus on preventing the personalization methods from generating high-quality corresponding images, often resulting in noticeable and unacceptable perturbations. They usually use $11/255$ [8, 15, 10] or $17/255$ [14] as perturbation size in their paper, which is generally unacceptable for owners of face photos, significantly limiting the usability of these methods in real-world applications. Therefore, we pose the following question in this work: How can we find a better trade-off between visibility of perturbations and protection effectiveness?
@@
  • We propose a visual-friendly concept protection framework to prioritize protecting essential information within images by a regional adversarial loss and loosen the optimization objective from finding the strongest perturbations within the perturbation size to the smallest feasible perturbations.
@@ II Related Work
Personalization of Diffusions Models. Personalization for specific concepts (attributions, styles, or objects) has been a long-standing goal in the image generation field. In text-to-image diffusion models, previous researchers have primarily concentrated on prompt learning and test-time tuning of pre-trained models to generate images based on specific target concepts using special language tokens. Textual Inversion adjusts text embeddings of a new pseudoword to describe the concept [6]. DreamBooth fine-tunes denoising networks to connect the novel concept and a less commonly used word token [7]. Based on that, more works [16, 17, 18] are proposed to improve controllability and flexibility in processing image visual concepts. In this paper, we have selected Textual Inversion and DreamBooth as the techniques used by the adversary due to their popularity and representativeness.
@@
Imperceptibility Adversarial Attack. Adversarial examples [19, 20, 21, 9, 22] are initially introduced by adding imperceptible noise to original data, fooling classifiers into misclassifying with high confidence. Recently, more and more researchers aim to improve the imperceptibility of adversarial examples, and they make use of a variety of tools such as perceptual color distance [23], low-frequency spatial constraints [24], hybrid attacks in frequency and spatial domain [25], and invertible neural networks [26], etc. But they are mainly aimed at discriminative tasks such as image classification, where important information in the image can be fed back relatively accurately by the gradient of the target model, whereas we target a diffusion-based generative model whose gradient is also for the whole image.
Adversarial Examples Against Unauthorized AI Generation. Unauthorized AI Generation seriously harms the safety of society, prompting researchers to explore multiple approaches for mitigating these risks. While passive defense methods focus on distinguishing between authentic and synthetic images [27, 28], adversarial attacks are a promising way against unauthorized AI generation [29, 30, 31, 32]. Ruiz et al. [29] are the first to adapt adversarial attack methods to undermine the generation of realistic deepfake images, thereby enhancing privacy security. Anti-Forgery [30] computes adversarial noise in the Lab color space to enhance the imperceptibility and robustness of various image transformations. Unlike others, DeepFake Disrupter [31] generates adversarial noise using an autoencoder instead of gradients and spoofs deepfake generators while remaining undetected by deepfake detectors. Zhu et al. [32] utilize adversarial examples to protect images while simultaneously recovering original image information from tampered images. In contrast to discriminative tasks, adversarial examples play a positive role in the image generation field.
With the rising popularity of diffusion models, previous works are difficult to adapt due to the obvious differences between diffusion and GAN or other generation models. Therefore, there is a growing interest in using adversarial examples to protect images from being maliciously exploited by diffusion models [12, 13, 10]. Photoguard [12] attacks the VAE encoder or whole model to prevent images from being maliciously edited by diffusion models at the test time. At the same time, AdvDM [13] employs adversarial examples to protect artists from having their styles imitated by textual inversion by maximizing the denoising loss of a pre-trained diffusion model. As an upgraded version of AdvDM, Mist [14] adds a texture loss that measures the distance between the encoder representation of the original image and the perturbed image in the image encoder of latent diffusion. Similar to them, GLAZE [15] proposes to measure the art style similarity between the perturbed image and the target style using a pre-trained style-transfer model, and adversarial examples can protect the art style-related attributes in the original image directly. Different from them, Anti-DreamBooth [8] protects privacy information from being learned by DreamBooth. Based on that, MetaCloak [10] uses a meta-learning strategy and data transformations to generate better effective and robust adversarial examples. Meanwhile, Xue et al. [11] introduce Score Distillation Sampling (SDS) loss into the perturbation optimization process to reduce computation cost and find maximize or minimize SDS loss both achieve competitive protection effects.
However, except for GLAZE, existing work protects all information for the whole image, which often requires significant noise, significantly affects the user experience, and may overemphasize the background, omitting the information that needs to be protected. In our work, we propose to prioritize limited noise to protect important information in the image, such as faces, and specific IPs. Unlike GLAZE which targets art style, we target specific semantic regions.
@@
Figure 2: The pipeline of VCPro. Initially, users can utilize SAM or alternative segmentation tools to create user-protective masks within images. Following this, the masks and the original images are simultaneously input into the protected image generation module, as described by Eq. (8). Within this module, we introduce a regional adversarial learning loss aimed at diminishing the visibility of protected images by precise protection and transform objectives to minimize perturbations while maintaining effective protection.
III Preliminaries
@@
Personalization based on Diffusion. Concept-driven personalization customizes the outputs of a generative model to align with a specific concept, such as attributes, style, objects, etc. Most personalization techniques are applied to the latent diffusion model (LDM, parameterized by $\theta$). A typical LDM consists of an image encoder $\mathcal{E}_{\theta}$, an image decoder $\mathcal{D}_{\theta}$, a condition encoder $\tau_{\theta}$, and a conditional denoising UNet $\varepsilon_{\theta}$. The training of LDMs contains diffusion and denoising processes, which both happen in the latent space. For an image $x$, $z_{0}=\mathcal{E}_{\theta}(x)$ is the latent code of $x$. The diffusion process randomly samples time $t$ and Gaussian noise $g$, and latent vector $z_{0}$ is added noise $g$ to $z_{t}$. Then the denoising UNet $\varepsilon_{\theta}$ is assigned to remove the noise added to the latent code $z_{t}$ conditioned by $\tau_{\theta}(y)$ in each timestep $t$, and we set the output of $\varepsilon_{\theta}$ as $g^{\prime}$. The training objective of this LDM is:
$\mathcal{L_{\theta}}:=\mathbb{E}_{z\sim\mathcal{E}(x),y,g\sim\mathcal{N}(0,1),t}\left[\left\|{g-\varepsilon_{\theta}\left(z_{t},t,\tau_{\theta}(y)\right)}\right\|_{2}^{2}\right],$ (1)
where $t$ is a timestep of the diffusion process, and $y$ is the given condition for generation.
Based on LDM, Textual Inversion, and DreamBooth are two main techniques for personalization. Textual Inversion learns new special tokens containing specific concepts to store new knowledge from images. In Textual Inversion, $y$ usually is a generic prompt, such as “a photo of sks [class noun]”, with $sks$ is a pseudo token, and “[class noun]” is the object type. They optimize the embedding of $sks$ directly by minimizing the objective specified in Eq. (1), which uses a small set of images representing the concept as training data.
@@
DreamBooth also binds new conceptual knowledge to specific words. Unlike Textual Inversion, it fine-tunes part of the parameters of LDMs on training images and class examples $x^{p}$, which is generated from original LDM with prior prompt $y^{p}$ to mitigate catastrophic forgetting.
@@
Protected Image for Diffusions. Protected images (adversarial examples) aim to add an imperceptible adversarial perturbation $\delta$ to the original image $x$ to make the full process of condition generation ineffective. The protected image $\mathcal{L_{\theta}}$ can be formalized as:
$\begin{gathered}x^{\prime}:=arg\max_{x^{\prime}}\mathcal{L_{\theta}}(x^{\prime},y),\\ \text{ s.t. }\quad\left\|x-x^{\prime}\right\|\leq\epsilon,\end{gathered}$ (4)
where $\epsilon$ is the perturbation size, also known as the perturbation budget, and limits the maximum range of adversarial perturbation. Adversarial attack transforms original images into special samples that are difficult to denoise. They enhance the optimization challenge of LDMs, thereby achieving image protection.
IV Visual-Friendly Concept Protection
IV-A Overview
Fig. 2 shows the pipeline of the proposed framework for visual-friendly concept protection. Accurately describing spatial positions through language can be challenging for users, but it can be precisely achieved using masks. By leveraging SAM [33] or other segmentation tools, users can generate a mask $m$ for important concepts within a given image $x$. The user-provided masks and original images are then collectively fed into the protected image generation module as described in Eq. (8). In this module, we propose a regional adversarial learning loss to reduce the visibility of protected images through precise protection and a Lagrangian multiplier-based solution to minimize perturbations while maintaining successful protection.
In this section, we start by formulating the regional adversarial learning framework for diffusion models in Sec. IV-B and then move on to the solution of the proposed optimization objectives in Sec. IV-C. The total learning process is shown in Alg. 1.
IV-B Formulation
@@
$\begin{gathered}x^{\prime}:=\arg\max_{x^{\prime}}\mathcal{L}^{\prime}_{\theta}\left(x^{\prime},y,m\right),\\ \text{ s.t. }\quad\left\|x-x^{\prime}\right\|\leq\epsilon,\\ \end{gathered}$ (5)
@@
$l_{mask}(z_{t},y,g,t,m):=$ $\underbrace{\left\|(g-\varepsilon_{\theta}\left(z_{t},t,\tau_{\theta}(y)\right))\odot m\right\|_{2}^{2}}_{\text{push term for protected regions}}$
$-\underbrace{\left\|(g-\varepsilon_{\theta}\left(z_{t},t,\tau_{\theta}(y)\right))\odot(1-m)\right\|_{2}^{2}}_{\text{pull term for non-protected regions}}.$
This loss operates through a balanced mechanism of opposing forces. The push term maximizes the distance between the predicted noise $\varepsilon_{\theta}$ and ground truth noise $g$ in masked regions ($m=1$), effectively disrupting the denoising process for protected concepts. Simultaneously, the pull term minimizes this distance in unmasked regions ($m=0$), preserving visual quality in non-protected areas.
During optimization, these components generate distinct gradient signals: push gradients divert predictions away from ground truth in protected regions, while pull gradients maintain accuracy elsewhere. This dual approach is crucial for effective protection—without the pull component, regions outside the mask cannot provide sufficient gradient feedback, significantly impairing the optimization process. Our ablation studies in Fig. 4 and Table II demonstrate that removing the pull component substantially degrades protection effectiveness.
@@
IV-C Lagrangian Multiplier-based Loose Solution
Based on Eq. (6), protected images will destroy the target area as much as possible. However, for the generation task, the output can still be recognized as synthetic as long as there are some clear signs of protection. Typical image classification attacking methods [20, 24, 26] also present similar viewpoints that minimize the impact on normal visual perception while maintaining the effectiveness of the attack. Hence, we propose a loosed optimization objective by attempting to find the minimal adversarial perturbation $\delta$ that can attack diffusion models successfully. For convenience, we set a heuristic method: if $\mathcal{L}^{\prime}_{\theta}>\alpha$, the attack is successful. The problem can be translated as:
@@ IV-C Lagrangian Multiplier-based Loose Solution
Considering that the human visual system is more sensitive to low-frequency regions [24], we use $D(\cdot)$ to limit perturbations into the high-frequency regions. Specifically, we use discrete wavelet transform (DWT) to transform images from the spatial domain to the frequency domain. DWT will decompose the image $x$ into one low-frequency and three high-frequency components, i.e., $x_{ll},x_{lh},x_{hl},x_{hh}$, and inverse DWT (IDWT) uses all four components to reconstruct the image.
@@
In this work, we drop the high-frequency components and reconstruct an image with only the low-frequency component as $\widetilde{x}=\phi(x)$, where $\phi(x)=L^{T}x_{ll}L=L^{T}(LxL^{T})L,\text{and }D(\cdot)=||\widetilde{x}-\widetilde{x^{\prime}}||_{2}^{2}$.
V Experiments
@@
V-A Experimental Settings
Datasets. We used the CelebA-HQ [34] and VGGFace2 [35] datasets, following Anti-DreamBooth [8] and MetaCloak [10], to validate the effectiveness of our method in privacy protection. CelebA-HQ is an enhanced version of the original CelebA dataset consisting of $30,000$ celebrity face images. VGGFace2 is a large-scale dataset with over $3.3$ million face images from $9,131$ unique identities. We randomly selected 50 identities from each dataset, ensuring that each had at least 15 images with a resolution exceeding 500 × 500 pixels. We also provide other non-face cases (artwork and landscape) in the paper.
Training details. The availability of Stable Diffusion pre-trained weights on Hugging Face  [^1] https://huggingface.co/ has significantly advanced research within the community. Here, our experiments primarily focus on the Stable Diffusion v1-4 version. For image pre-processing, we center crop and resize images to a resolution $512\times 512$. During protected image generation, we use $720$ iterations, step size $\alpha$ is $1/255$, and adversarial perturbation size $\epsilon$ is $8/255$. We test the protection performance on two types of SOTA personalization methods, i.e., Textual Inversion (TI) [6], and DreamBooth (DB) [7]. During the training of the Textual Inversion, the constant learning rate is set to $5\times 10^{-4}$ with $3000$ optimization steps and batch size $1$. In DreamBooth, the constant learning rate is set to $5\times 10^{-7}$ with $1000$ optimization steps and batch size is $2$. We use a guidance scale of $7.5$ and $50$ denoising steps at the generation phase with a testing prompt. We set $c=0.1$ and $\alpha=0.5$ in practice. All experiments are complete on NVIDIA A100 GPU 40GB.
Baselines. We compare our method with five SOTA baselines, i.e., Mist [13], Anti-DreamBooth (referred to as Anti-DB) [8], PhotoGuard [12], SDS(-) [11] and MetaCloak [10]. In implementations, we set $\epsilon=8/255$ as the $\ell_{\infty}$ perturbation size, $\alpha=1/255$ as the step size for fair comparison.
  • Mist aims to maximize the training loss of Latent Diffusion Models (LDMs) while minimizing a textual loss between the protected image and a dummy image in the image encoder layer during protected image generation. In our implementation, following the original paper [14], we set the loss weight for the textual loss to $1\times 10^{-4}$ and run 720 iterations during protected image generation.
  • PhotoGuard targets diffusion-based inpainting/editing methods and offers two variants: one that attacks the VAE-encoder of LDMs, and another that attacks the final image outputs. We select the VAE-encoder variant due to its superior visual quality and set the training iterations to 200, as recommended in the original paper [12].
  • SDS(-) minimizes the SDS loss [36] to optimize adversarial data, achieving the best visual quality among all versions presented in its original paper [11]. We set the training iterations to 100, following [11].
  • MetaCloak utilizes meta-learning to enhance the robustness of adversarial data against various transformations. We set the number of surrogate models to 5, the unrolling number to 1, the sample batch size to 1, and the training iterations to 4000, based on the recommendations in [10].

### v3

@@ Abstract
Personalized concept generation by tuning diffusion models with a few images raises potential legal and ethical concerns regarding privacy and intellectual property rights. Researchers attempt to prevent malicious personalization using adversarial perturbations. However, previous efforts have mainly focused on the effectiveness of protection while neglecting the visibility of perturbations. They utilize global adversarial perturbations, which introduce noticeable alterations to original images and significantly degrade visual quality. In this work, we propose the Visual-Friendly Concept Protection (VCPro) framework, which prioritizes the protection of key concepts chosen by the image owner through adversarial perturbations with lower perceptibility. To ensure these perturbations are as inconspicuous as possible, we introduce a relaxed optimization objective to identify the least perceptible yet effective adversarial perturbations, solved using the Lagrangian multiplier method. Qualitative and quantitative experiments validate that VCPro achieves a better trade-off between the visibility of perturbations and protection effectiveness, effectively prioritizing the protection of target concepts in images with less perceptible perturbations.
Introduction
@@
With the advent of popular image generative models (ho2020denoising; rombach2022high; song2020denoising) such as Stable Diffusion (rombach2022high) and GPT-4o (achiam2023gpt), people lacking expertise in drawing or photography can effortlessly create realistic or artistic works using simple textual descriptions. However, the success of these models has raised significant concerns about privacy, intellectual property rights, and various legal and ethical issues (luo2024exploring). For example, an adversary could easily generate sensitive specific concepts, such as personal fake images, or imitate renowned artworks for commercial purposes, using only a few reference images and concept personalization techniques such as Textual Inversion (gal2022image) or DreamBooth (ruiz2023dreambooth).
@@
Figure 1: The first row shows protected images from Anti-DreamBooth (le_etal2023antidreambooth), Mist (zhang2022towards), MetaCloak (liu2024metacloak), SDS(-) (xue2024toward), PhotoGuard (salman2023raising), and our VCPro. Second and third rows show magnified regions and Textual Inversion results. At $\epsilon$=8/255, our method better balances protection effectiveness and visual quality (higher PSNR/SSIM).
@@
One prevailing direction to mitigate these potential risks is to leverage adversarial attack techniques, transforming original images into adversarial examples, termed “protected images”. These protected images can misguide the personalized diffusion model and deceive its generation process, which resists malicious editing or personalization (salman2023raising; liang2023adversarial; liang2023mist; le_etal2023antidreambooth; shan2023glaze; liu2024metacloak). For instance, AdvDM (liang2023adversarial) employs adversarial attacks in the inversion stage of Stable Diffusion to safeguard against malicious imitation of the style of a specific artist. Later, its updated version, Mist (liang2023mist), enhances the protection efficacy of protected images by adding a textual loss, extending its application from Textual Inversion to DreamBooth. Unlike AdvDM and Mist, which focus on art style protection, Anti-DreamBooth (le_etal2023antidreambooth) undermines DreamBooth model generation quality to enhance privacy by adding adversarial perturbations to human face images before posting online. Furthermore, MetaCloak (liu2024metacloak) leverages a meta-learning strategy and data transformations to generate more effective and robust protected images against DreamBooth.
@@
However, these methods primarily focus on preventing the personalization methods from generating high-quality corresponding images, often resulting in noticeable and unacceptable perturbations. They usually use $11/255$ (le_etal2023antidreambooth; shan2023glaze; liu2024metacloak) or $17/255$ (liang2023mist) as perturbation size in their paper, which is generally unacceptable for owners of face photos, significantly limiting the usability of these methods in real-world applications. Therefore, we pose the following question in this work: How can we find a better trade-off between the visibility of perturbations and protection effectiveness?
@@
  • We propose a visual-friendly concept protection framework that uses regional adversarial loss to protect essential image information. Considering human sensitivity, we measure perturbation perceptibility in the frequency domain and optimize for the smallest feasible perturbations rather than the strongest ones within size constraints.
@@
Personalization of Diffusions Models. Personalization for specific concepts (attributions, styles, or objects) has been a long-standing goal in the image generation field. In text-to-image diffusion models, previous researchers have primarily concentrated on prompt learning and test-time tuning of pre-trained models to generate images based on specific target concepts using special language tokens. Textual Inversion adjusts text embeddings of a new pseudoword to describe the concept (gal2022image). DreamBooth fine-tunes denoising networks to connect the novel concept and a less commonly used word token (ruiz2023dreambooth). Based on that, more works (voynov2023p+; zhang2023prospect; kumari2023multi) are proposed to improve controllability and flexibility in processing image visual concepts. In this paper, we have selected Textual Inversion and DreamBooth as the techniques used by the adversary due to their popularity and representativeness.
@@
Imperceptibility Adversarial Attack. Adversarial examples (szegedy2013intriguing; carlini2017towards; duan2021advdrop; mi2023adversarial) are initially introduced by adding imperceptible noise to original data, fooling classifiers into misclassifying with high confidence. Recently, more and more researchers aim to improve the imperceptibility of adversarial examples, and they make use of a variety of tools such as perceptual color distance (zhao2021invertible), low-frequency spatial constraints (luo2022frequency), hybrid attacks in frequency and spatial domain (jia2022exploring), and invertible neural networks (chen2023imperceptible), etc. But they are mainly aimed at discriminative tasks such as image classification, where important information in the image can be fed back relatively accurately by the gradient of the target model, whereas we target a diffusion-based generative model whose gradient is also for the whole image.
Adversarial Examples Against Unauthorized Diffusion Generation. Unauthorized AI generation poses significant safety risks, driving research into mitigation approaches. While passive defenses focus on detecting synthetic images (wu2022robust; li2017localization), adversarial attacks offer promising protection against unauthorized generation (ruiz2020disrupting; wanganti; wang2022deepfake; zhu2023information). Recent works target diffusion models specifically. Photoguard (salman2023raising) attacks VAE encoders to prevent malicious editing, while AdvDM (liang2023adversarial) protects artistic styles by maximizing denoising loss. Mist (liang2023mist) enhances AdvDM with texture loss in latent space. GLAZE (shan2023glaze) measures art style similarity using pre-trained style-transfer models. Anti-DreamBooth (le_etal2023antidreambooth) protects privacy from DreamBooth learning, with MetaCloak (liu2024metacloak) improving robustness via meta-learning. Xue et al. (xue2024toward) introduce Score Distillation Sampling (SDS) loss to reduce computational costs. However, except for GLAZE, existing methods protect entire images with significant noise, degrading user experience and potentially overemphasizing backgrounds while missing critical information. We propose prioritizing limited noise to protect important semantic regions like faces and specific IPs, unlike GLAZE which targets art style.
@@
Figure 2: VCPro pipeline. Users create protective masks via SAM or other tools. The masks and images are fed into our protection module (Eq. 8), which uses regional adversarial learning to minimize perturbation visibility while maintaining protection.
Preliminaries
@@
Personalization based on Diffusion. Concept-driven personalization customizes generative model outputs to align with specific concepts. Most techniques apply to latent diffusion models (LDMs, parameterized by $\theta$) consisting of image encoder $\mathcal{E}_{\theta}$, decoder $\mathcal{D}_{\theta}$, condition encoder $\tau_{\theta}$, and denoising UNet $\varepsilon_{\theta}$. For image $x$ with latent code $z_{0}=\mathcal{E}_{\theta}(x)$, the training objective is:
$\mathcal{L_{\theta}}:=\mathbb{E}_{z\sim\mathcal{E}(x),y,g\sim\mathcal{N}(0,1),t}\left[\left\|{g-\varepsilon_{\theta}\left(z_{t},t,\tau_{\theta}(y)\right)}\right\|_{2}^{2}\right].$ (1)
Textual Inversion learns embeddings $v$ for pseudo tokens skssks in prompts like “a photo of skssks [class noun]” by optimizing:
@@
DreamBooth fine-tunes LDM parameters on training images and class examples $x^{p}$ to prevent catastrophic forgetting:
@@
Protected Image for Diffusions. Protected images add imperceptible adversarial perturbations $\delta$ to original images $x$ to disrupt personalization. The protected image $x^{\prime}$ is formalized as:
$\begin{gathered}x^{\prime}:=\arg\max_{x^{\prime}}\mathcal{L_{\theta}}(x^{\prime},y),\\ \text{s.t.}\quad\left\|x-x^{\prime}\right\|\leq\epsilon,\end{gathered}$ (4)
where $\epsilon$ limits the perturbation budget. These adversarial attacks create samples difficult to denoise, enhancing LDM optimization challenges for image protection.
Figure 3: Qualitative comparison ($\epsilon=8/255$, SD v1-4). Rows show original and protected images from baseline methods and ours, with TI/DB results. Perturbations visualized in black-yellow. Please zoom in.
Visual-Friendly Concept Protection
Overview
Fig. 2 shows the pipeline of the proposed framework for visual-friendly concept protection. Accurately describing spatial positions through language can be challenging for users, but it can be precisely achieved using masks. By leveraging SAM (Kirillov_2023_ICCV) or other segmentation tools, users can generate a mask $m$ for important concepts within a given image $x$. The user-provided masks and original images are then collectively fed into the protected image generation module as described in Eq. (8). In this module, we propose a regional adversarial learning loss to reduce the visibility of protected images through precise protection and a Lagrangian multiplier-based solution to minimize perturbations while maintaining successful protection.
In this section, we start by formulating the regional adversarial learning framework for diffusion models in Sec. Formulation and then move on to the solution of the proposed optimization objectives in Sec. Lagrangian Multiplier-based Loose Solution. The total learning process is shown in Alg. 1.
Formulation
@@
[…]
@@
$l_{mask}:=$ $\left\|\Delta\odot m\right\|_{2}^{2}-\left\|\Delta\odot(1-m)\right\|_{2}^{2},$
where $\Delta=g-\varepsilon_{\theta}(z_{t},t,\tau_{\theta}(y))$. This loss operates through a balanced mechanism of opposing forces. The push term maximizes the distance between the predicted noise $\varepsilon_{\theta}$ and ground truth noise $g$ in masked regions ($m=1$), effectively disrupting the denoising process for protected concepts. Simultaneously, the pull term minimizes this distance in unmasked regions ($m=0$), preserving visual quality in non-protected areas.
During optimization, these components generate distinct gradient signals: push gradients divert predictions away from ground truth in protected regions, while pull gradients maintain accuracy elsewhere. This dual approach is crucial for effective protection—without the pull component, regions outside the mask cannot provide sufficient gradient feedback, significantly impairing the optimization process. Our ablation studies in Table 3 demonstrate that removing the pull component substantially degrades protection effectiveness.
@@
Lagrangian Multiplier-based Loose Solution
Based on Eq. (6), protected images will destroy the target area as much as possible. However, for the generation task, the output can still be recognized as synthetic as long as there are some clear signs of protection. Typical image classification attacking methods (carlini2017towards; luo2022frequency; chen2023imperceptible) also present similar viewpoints that minimize the impact on normal visual perception while maintaining the effectiveness of the attack. Hence, we propose a loosed optimization objective by attempting to find the minimal adversarial perturbation $\delta$ that can attack diffusion models successfully. For convenience, we set a heuristic method: if $\mathcal{L}^{\prime}_{\theta}>\alpha$, the attack is successful. The problem can be translated as:
@@
Considering that the human visual system is more sensitive to low-frequency regions (luo2022frequency), we use $D(\cdot)$ to limit perturbations into the high-frequency regions. Specifically, we use discrete wavelet transform (DWT) to transform images from the spatial domain to the frequency domain. DWT will decompose the image $x$ into one low-frequency and three high-frequency components, i.e., $x_{ll},x_{lh},x_{hl},x_{hh}$, and inverse DWT (IDWT) uses all four components to reconstruct the image.
@@
[…]
Dataset Method Visibility TI DB Human Eval
[…]
VGGFace2
No Defense 3.61±0.98 0.24±0.34 3.95±1.11 0.38±0.25 64/25/11
Mist 105.9±9.8 0.83±0.03 32.5±1.1 1.24±0.41 0.11±0.29 1.02±0.11 0.25±0.28 4/6/90 98.3 97.2
Anti-DB 96.2±10.4 0.82±0.03 32.4±1.1 1.18±0.36 0.05±0.29 1.03±0.14 0.26±0.28 3/9/88 97.4 100.0
PhotoGuard 62.1±8.2 0.84±0.03 33.2±1.2 1.50±0.54 0.13±0.34 1.26±0.30 0.27±0.31 13/19/68 98.7 84.0
SDS(-) 47.5±6.8 0.81±0.03 32.9±0.4 2.09±0.82 0.07±0.37 2.74±0.97 0.33±0.26 13/16/71 94.6 77.1
MetaCloak 204.3±54.5 0.82±0.02 32.1±0.6 1.86±0.62 0.19±0.33 1.33±0.41 0.25±0.32 2/10/88 91.0 90.4
VCPro 27.0±4.4 0.90±0.03 35.3±1.8 2.31±0.78 0.03±0.32 2.06±0.70 0.21±0.34 98.6 97.4
CelebA-HQ
No Defense 4.40±0.86 0.41±0.22 4.84±0.40 0.63±0.13 77/19/4
Mist 78.3±11.8 0.86±0.03 33.8±0.2 1.45±0.58 0.26±0.26 1.04±0.11 0.44±0.21 5/4/91 96.3 100.0
Anti-DB 78.5±11.7 0.86±0.03 33.7±0.2 1.81±1.20 0.28±0.25 1.04±0.15 0.44±0.21 0/4/96 97.0 96.5
PhotoGuard 45.2±7.2 0.88±0.02 35.1±0.3 1.81±0.59 0.30±0.24 1.23±0.35 0.48±0.17 2/22/76 94.3 94.3
SDS(-) 34.5±6.2 0.82±0.03 33.0±0.3 2.29±0.70 0.09±0.28 2.92±0.74 0.55±0.15 9/11/80 98.6 84.3
MetaCloak 161.9±26.9 0.86±0.03 33.1±0.2 1.47±0.56 0.25±0.26 1.59±0.53 0.44±0.16 0/4/96 96.0 95.5
VCPro 16.2±2.9 0.95±0.01 39.3±0.4 2.61±1.00 0.09±0.28 2.62±0.95 0.44±0.19 100.0 95.7
Table 1: Comparison of protection methods on VGGFace2 and CelebA-HQ with $\epsilon=8/255$. Metrics include visibility (FID, SSIM, PSNR), quality degradation (LIQE, CLIP-FACE), and protection efficacy (TI/DB: % synthetic images detected). Human evaluation visibility shows Lose/Tie/Win rates for VCPro vs. baselines. $\uparrow$$\downarrow$ indicates higher/lower is better. Bold: best performance.
Experiments
@@
Experimental Settings
We validate our method on CelebA-HQ (karras2017progressive) and VGGFace2 (cao2018vggface2) datasets, randomly selecting 50 identities from each with at least 15 images exceeding 500×500 resolution, following Anti-DreamBooth (le_etal2023antidreambooth) and MetaCloak (liu2024metacloak). Using Stable Diffusion v1-4 with 512×512 resolution, we generate protected images with 720 iterations, step size $\alpha=1/255$, and perturbation size $\epsilon=8/255$. We test protection performance against Textual Inversion (gal2022image) (learning rate $5\times 10^{-4}$, 3000 steps) and DreamBooth (ruiz2023dreambooth) (learning rate $5\times 10^{-7}$, 1000 steps), comparing with five SOTA baselines: Mist (liang2023adversarial), Anti-DreamBooth (le_etal2023antidreambooth), PhotoGuard (salman2023raising), SDS(-) (xue2024toward), and MetaCloak (liu2024metacloak). We evaluate visual perception using FID (heusel2017gans), SSIM (wang2004image), and PSNR between protected and original images, and assess protection effectiveness using LIQE (zhang2023blind) for full image quality and CLIP-FACE (liu2024metacloak) for face regional quality evaluation. LIQE measures image quality on a five-point scale: $c\in C=\{1,2,3,4,5\}=\{\text{``bad''},\text{``poor''},\text{``fair''},\text{``good''},\text{``perfect''}\}$. CLIP-FACE is based on CLIP-IQA for visual quality by considering additional class information. All experiments are conducted on NVIDIA A100 GPU 40GB with parameters $c=0.1$ and $\alpha=0.5$. We provided detailed experimental settings in Supplementary Materials.
Method Visibility TI DB